OpenStudy (anonymous):

evaluate the following expression: 3(1,-1)-(4,-2) a- (3,2) b- (0,4) c- (-1,-1) d- (4,-1)

OpenStudy (goformit100):

should i solve it for you or give you a hint

OpenStudy (anonymous):

If you could solve it that would be great, Im not sure how to do it.

OpenStudy (goformit100):

3(1,-1)-(4,-2) =(3,-3)-(4,-2) =(-1,-1) Ans.

OpenStudy (anonymous):

Thanks(:

OpenStudy (goformit100):

welcome!!!